### v1

# Dolphin: A Programmable Framework for Scalable Neurosymbolic Learning

Abstract
Neurosymbolic learning has emerged as a promising paradigm to incorporate symbolic reasoning into deep learning models. However, existing frameworks are limited in scalability with respect to both the training data and the complexity of symbolic programs. We propose Dolphin, a framework to scale neurosymbolic learning at a fundamental level by mapping both forward chaining and backward gradient propagation in symbolic programs to vectorized computations. For this purpose, Dolphin introduces a set of abstractions and primitives built directly on top of a high-performance deep learning framework like PyTorch, effectively enabling symbolic programs to be written as PyTorch modules. It thereby enables neurosymbolic programs to be written in a language like Python that is familiar to developers and compile them to computation graphs that are amenable to end-to-end differentiation on GPUs. We evaluate Dolphin on a suite of 13 benchmarks across 5 neurosymbolic tasks that combine deep learning models for text, image, or video processing with symbolic programs that involve multi-hop reasoning, recursion, and even black-box functions like Python eval(). Dolphin only takes $0.33$%-$37.17$% of the time (and $2.77$% on average) to train these models on the largest input per task compared to baselines Scallop, ISED, and IndeCateR+, which time out on most of these inputs. Models written in Dolphin also achieve state-of-the-art accuracies even on the largest benchmarks.
1 Introduction
Deep learning has achieved remarkable success on a wide range of tasks such as image classification, speech recognition, and natural language processing. With the recent emergence of foundation models like GPT-4 and CLIP, deep learning models are now increasingly being used in more challenging tasks. While such models work well for prediction and generation tasks, they are limited in their ability to perform reasoning required for tasks involving structure, logic, and planning, where symbolic reasoning traditionally excels (Kambhampati et al., 2024). Neurosymbolic programming (Chaudhuri et al., 2021) has emerged as a promising paradigm to incorporate symbolic reasoning into deep learning models, providing the best of both worlds.
Various frameworks have been developed to improve the programmability and accessibility of neurosymbolic applications (Manhaeve et al., 2018; Li et al., 2023; Solko-Breslin et al., 2024). These frameworks support complex symbolic reasoning features like recursion and black-box functions, implement efficient differentiable reasoning algorithms, and provide bindings for deep learning frameworks like PyTorch. Despite significant strides, however, these frameworks incur significant overhead when training neurosymbolic programs.
Consider a typical workflow of such a framework depicted in Figure 1(a). In this case, we have a supervised learning task with labeled data $(x,y)$, a neural network $M_{\theta}$ that processes input $x$, and a symbolic program $P_{\tt symbolic}$ that takes the network’s output $r$ and produces final output $y$. Existing frameworks, such as Scallop (Li et al., 2023), execute the neural model on GPU but use a separate CPU-based backend (implemented in Rust in Scallop’s case) for the symbolic program. Moreover, such systems suffer from slow inter-process data transfers between the two environments.
(a) Typical neurosymbolic framework (e.g. Scallop).
(b) Our neurosymbolic framework Dolphin.
Figure 1: Comparison of system architectures of neurosymbolic frameworks.
Together, these issues hinder the scalability of neurosymbolic learning with respect to problem complexity and data complexity. First, the symbolic computation engine must derive a set of all possible results and their associated probabilities in a manner that is differentiable with respect to the network’s parameters $\theta$. As the complexity of the symbolic program increases, the number of possible results and their associated weights also grows exponentially, leading to a combinatorial explosion in the number of required computations. However, the symbolic computations are discrete and not easily parallelizable on modern hardware like GPUs. Second, larger datasets also compound the computational cost of neurosymbolic learning. Deep learning typically addresses this challenge by batching computations across multiple data samples. However, in neurosymbolic learning, the computations may differ across data samples, making it difficult to batch them effectively.
To address these challenges, we need to fundamentally rethink the design of a neurosymbolic framework. One approach is to develop specialized and low-level primitives that scale specific benchmarks but make it time-intensive for developers to write neurosymbolic programs tailored to particular tasks. Alternatively, providing high-level primitives—such as a logic programming language like Scallop (Li et al., 2023) or DeepProbLog (Manhaeve et al., 2018)—simplifies the development of symbolic programs but limits the fine-grained control needed to scale specific applications. Finally, to truly democratize neurosymbolic programming, it is crucial to develop a framework that seamlessly integrates into the everyday deep learning workflows that developers already use.
In this work, we propose Dolphin, a novel framework for scalable neurosymbolic learning. In Dolphin, we build three key components that effectively tackle the scalability and programmability challenges described above. First, we develop a general symbolic representation that efficiently captures the relationships between neural network outputs and associated discrete symbols. Second, we introduce a set of primitives to map forward chaining in symbolic programs to vectorized computations over these representations. Third, we develop a set of vectorized provenance semirings (Green et al., 2007) that are easily pluggable into Dolphin and enable to efficiently compute symbolic gradients. As illustrated in Figure 1(b), these components together allow Dolphin to build a computation graph that spans both symbolic and neural operations, is highly parallelizable, and end-to-end differentiable on GPUs. Finally, Dolphin is implemented as a library that is integrated with PyTorch, allowing users to easily incorporate it into their existing deep learning pipelines.
We evaluate Dolphin on a diverse set of neurosymbolic tasks that involve text, image, video, and multi-modal data, and use rich reasoning features such as recursion and black-box Python functions. Neurosymbolic programs written using Dolphin only require $0.33$%-$37.17$% ($2.77$% on average) of the time to train compared to state-of-the-art baselines including differentiable reasoning frameworks like Scallop, and sampling-based frameworks like ISED and IndeCateR+ while maintaining similar levels of accuracy. We also observe that Dolphin efficiently scales to more complex benchmarks and larger datasets whereas the baselines either time out after 10 hours or fail to converge.
We make the following contributions in this work:
  • We propose Dolphin, a novel neurosymbolic programming framework for end-to-end differentiable symbolic reasoning in a scalable manner.
  • We develop novel abstractions to represent symbolic and neural computations and introduce vectorized primitives for neurosymbolic programs.
  • We develop vectorized provenances that can be plugged into Dolphin for efficient computation of symbolic gradients on parallelizable hardware such as GPUs.
  • We evaluate Dolphin on a diverse range of challenging neurosymbolic tasks across different domains and show that it effectively scales with increasing problem complexity and dataset size.
2 Overview
We illustrate our approach using the MNIST Sum-$N$ task from (De Smet et al., 2024). The goal is to train a model that takes as input $N$ images of MNIST digits and returns the sum of the digits represented by the images. During learning, supervision is provided only on the sum instead of the labels of the digits. The difficulty of the problem scales exponentially as there are $10^{N}$ states in the input space. Further, there are only $19N+$ 1 possible labels, resulting in very sparse supervision.
1class MNISTNet(nn.Module):
2 def __init__(self):
3 super(MNISTNet, self).__init__()
4 ...
5
6class SumNNet(nn.Module):
7 def __init__(self):
8 super(SumNNet, self).__init__()
9 self.CNN = MNISTNet()
10
11 def forward(self, imgs):
12 digits = range(10)
13 D_res = Distribution(self.CNN(imgs[0]), digits)
14 for i in range(1, len(imgs)):
15 D_i = Distribution(self.CNN(imgs[i]), digits)
16 D_res = apply(D_res, D_i, lambda x,y: x + y)
17 l_res = get_logits(D_res)
[…]
34
35
36
37
38
39
40
41
42
[…]
49
50In order to support training, the symbolic program must reason over all the outputs of the CNN, and return probability distributions over all the possible results (\$0\$ to \$9N\$).
[…]
52The batched nature of data in machine learning further complicates these calculations since the probabilities of symbols can be different across samples within the same batch.
53As a result, writing neurosymbolic programs in native PyTorch is tedious even for simple tasks.
54
[…]
59
[…]
65This is repeated iteratively until all the outputs of the CNN are summed appropriately.
[…]
67
[…]
89
[…]
3 The Dolphin Framework
3.1 Dolphin Core Design Principles
We based Dolphin’s framework design on the following core principles:
  • Flexible programmability: The framework should allow developers to write neurosymbolic applications in Python with minimal effort, providing intuitive primitives that seamlessly integrate with Python’s rich and expressive language features.
  • End-to-end differentiability on GPUs: The framework should allow any neurosymbolic program to be end-to-end differentiable on GPUs irrespective of the task characteristics.
  • Scalable: The framework should easily scale with greater problem and data complexity.
  • Tunable: Similar to hyperparameters in deep learning, the framework should provide a simple interface for developers to choose provenances (and their configurations) or define new ones.
Next, we show how we realize these principles by describing the key components of Dolphin.
3.2 The Dolphin Syntax
Dolphin provides a programming interface that developers can use to express symbolic programs in a Pythonic manner. Dolphin maps each operation of the symbolic program to PyTorch which enables end-to-end GPU-accelerated differentiable reasoning. Figure 3 presents Dolphin’s programming interface including the symbolic abstractions and operations over them.
$\begin{array}[]{rcrl}\text{Symbol}&::&s&\in\ \ S\ \ \text{ ({objects})}\\ \text{Tag}&::&t&\in\ \ T\ \ \text{ ({tensors})}\\ \text{Distribution}&::&D&\in\ \ \mathbb{D}=S\rightarrow T\end{array}$
$\begin{array}[]{rcl}\textsc{Apply}&:&\mathbb{D}^{K}\times(S^{K}\rightarrow S)\rightarrow\mathbb{D}\\ \textsc{Filter}&:&\mathbb{D}\times(S\rightarrow\mathbb{B})\rightarrow\mathbb{D}\\ \textsc{ApplyIf}&:&\mathbb{D}^{K}\times(S^{K}\rightarrow S)\times(S^{K}\rightarrow\mathbb{B})\rightarrow\mathbb{D}\\ \textsc{Union}&:&\mathbb{D}\times\mathbb{D}\rightarrow\mathbb{D}\\ \textsc{GetProbs}&:&\mathbb{D}\rightarrow[0,1]^{N}\end{array}$
Figure 3: Formal definition of Dolphin’s programming abstractions and primitives.
3.2.1 Abstractions
The three main abstractions provided by Dolphin for expressing differentiable symbolic programs are shown on the left of Figure 3. Symbols $S$ represent symbolic entities relevant to the program. These entities can be any Pythonic object, such as hand-written digits in MNIST-SumN or coordinates of points in PathFinder. Tags $T$ are tensors that represent relative likelihoods. Typically, tags for symbols are derived from the outputs of machine learning models, such as the probability distribution over digits produced by the CNN classifier in MNIST-SumN. Finally, Distribution $D$ represents the likelihood of an input being classified as one of the pre-defined symbols.
Distributions serve as the fundamental datatype of a Dolphin program and act as its main interface with a machine learning model. For instance, when the developer instantiates a Distribution object, such as in the following code snippet from Figure LABEL:fig:mnist_sum_n_code:
$\texttt{D}_{\text{res}}$ = Distribution(self.CNN(imgs[0]), digits)
the output of the CNN model is directly passed to the Distribution object, effectively acting as an input to the symbolic program. The Distribution object itself, as shown in Figure LABEL:fig:mnist_sum_breakdown, contains batches of tags extracted from the model outputs, and maintains the set of corresponding symbols.
To enable such a seamless integration between the PyTorch model and the symbolic program, Distributions are designed to operate directly over PyTorch tensors. This has two main advantages. First, it preserves the gradients of the model output throughout the symbolic program, enabling end-to-end differentiability. Second, it allows Dolphin to perform operations over an entire batch of tags, leveraging the vectorized operations provided by PyTorch. Dolphin can thus operate efficiently on specialized hardware like GPUs, allowing the symbolic program to scale effectively.
3.2.2 Operations
Figure 3 shows the five operations supported by Dolphin that developers can use to manipulate Distributions and express complex symbolic programs. We now expand on these operations.
Apply.
This is the primary operation that developers can use to manipulate Distributions. It takes as inputs $K$ Distributions, where $K\geq 1$, along with a function $f$ of the same arity. This function defines operations over the symbols of $K$ distributions. Apply then computes the results of $f$ over all possible combinations of arguments sourced from the symbols of the Distributions as well as their associated tags, and returns a new Distribution with these results and tags.
This operation occurs in two stages akin to the popular map-reduce pattern. In the map stage, Apply computes the results of $f$ over the symbols of the input Distributions and their associated tags:
$R\ =\ \{\ (f(s_{1},s_{2},\ldots,s_{k}),(t_{1}\otimes t_{2}\otimes\ldots\otimes t_{k}))\ |\ D_{i}(s_{i})=t_{i},i=1,\ldots,k\ \}$ (1)
Here, the tag of each result symbol $f(s_{1},s_{2},\ldots,s_{k})$ is the conjunction $\otimes$ of the tags $(t_{1},t_{2},\ldots,t_{k})$ of the input symbols it was derived from.
Note that the function $f$ can be any user-defined Python function, including complex functions like regex parsing, image processing, or Python’s eval(). As a result, $f$ may be a many-to-one function and the tags shared by a resulting symbol must be aggregated to form the final tags of the output Distribution. We, therefore, shuffle the results from the map stage to compute a function $M$ from each symbol to tags from $R$ associated with it:
$M=\lambda\ s\ .\ \{\ t\ |\ (s,t)\in R\ \}$ (2)
We then proceed to the reduce stage, where we aggregate the tags of each symbol in $M$ using disjunction $\oplus$ to produce the final Distribution $D_{\text{res}}$:
$D_{\text{res}}=\lambda\ s\ .\ \bigoplus\ \{\ t\ |\ t\in M(s)\ \}$ (3)
Since the tags here are PyTorch tensors representing probabilities, the implementations of the conjunction and disjunction operations are dictated by the underlying provenance used by the program. A more detailed explanation of the provenances is provided in Section 3.3.
Filter. The Filter operation is used to filter out symbols from a Distribution based on some condition. It takes in a single Distribution, along with a user-defined function that returns a boolean value, which acts as the condition. This operation then returns a new Distribution that contains only the symbols that satisfy the condition, along with their tags.
ApplyIf. The ApplyIf operation is a conditional version of Apply. It takes in $K$ Distributions and two functions $f_{\textit{apply}}$ and $f_{\textit{cond}}$ of the same arity. For each combination of symbols from the $K$ Distributions, ApplyIf computes the results of $f_{\textit{apply}}$ and its associated tags only if the condition $f_{\textit{cond}}$ is satisfied over that combination of symbols. The operation then returns a new Distribution with these results and tags.
Union. The Union operation is used to combine two Distributions. It takes in two Distributions and returns a new Distribution that contains the union of the symbols from the two input Distributions, along with their tags. Any symbols common to both input Distributions have their tags merged via a disjunction operation. Union is especially useful when writing recursive programs in Dolphin that require combining the results of multiple recursive calls.
GetProbs. The GetProbs operation is used to extract the probabilities from the tags of a Distribution. This is used mainly once the symbolic program has been executed to extract the final probabilities of the symbols in the output Distribution. These probabilities can then be used to compute the loss function for training the neural components of the neurosymbolic model. The actual extraction of the probabilities from the tags depends on the specific provenance used in the program, which we discuss next.
3.3 Dolphin Provenances
As discussed earlier, each symbol in a distribution is associated with a batch of one or more tags. The Dolphin primitives define how to manipulate certain tags. For instance, Equations (1) and (3) specify the tags to be conjuncted or disjuncted. We now define the semantics of these operations.
The goal of such operations is to approximate the probabilities of the symbols in the output distribution as accurately as possible. This is achieved by using a mathematical framework called provenance semirings (Green et al., 2007). Provenance semirings provide generalized algebraic structure to propagate tags when computing over tagged data. In the case of Dolphin distributions, we can view the tags as representing the probabilities, and the data as the distribution’s symbols.
Designing and implementing provenances can be challenging since they must be accurate enough to capture the semantics of the symbolic program, while at the same time being coarse enough to maintain computational feasibility. Furthermore, the provenances must be differentiable to enable end-to-end training for neurosymbolic tasks. While neurosymbolic frameworks like Scallop (Li et al., 2023) implement differentiable provenances, they are not designed to leverage hardware accelerations or batched optimizations due to the CPU-bound nature of their implementations. We thus design vectorized provenances in Dolphin that are differentiable and enable GPU computations.
We simplify the definition of provenances from Scallop as a 5-tuple: $(T,\mathbf{0},\mathbf{1},\otimes,\oplus)$. Here, $T$ is the tag space, $\otimes:T\times T\to T$ is the conjunction operator with identity $\mathbf{0}$, and $\oplus:T\times T\to T$ is the disjunction operator with identity $\mathbf{1}$. We then implement two differentiable provenances in Dolphin: Differentiable Add-Mult Probabilities (DAMP) and Differentiable Top-K Proofs (DTKP). Table 1 summarizes the operations of these provenances.
Provenance Domain $\mathbf{0}$ $\mathbf{1}$ $t\oplus t^{\prime}$ $t\otimes t^{\prime}$
DAMP $[0,1]$ 0 1 $\text{clamp}_{0}^{1}(t+t^{\prime})$ $t\cdot t^{\prime}$
DTKP-AM $[0,1]\cup\{\infty,-\infty\}$ $\mathbf{\hat{0}}_{ij}=\begin{cases}\infty&i=1\\ -\infty&i>1\end{cases}$ $\mathbf{\hat{1}}_{ij}=\infty$ $\text{top}_{k}(\text{cat}(t,t^{\prime}))$ $\text{top}_{k}([\min(|t_{i}|,|t^{\prime}_{j}|)\mid(t_{i},t^{\prime}_{j})\in t\times t^{\prime}])$
Table 1: Dolphin provenances implemented in PyTorch.
Differentiable Add-Mult Probabilities. Differentiable Add-Mult Probabilities (DAMP) is a popular technique that uses the probability space as its tag space: $T=[0,1]$. Its conjunction operation $\otimes$ is defined as the addition of probabilities, clamped at $\mathbf{1}$, and its disjunction operation $\oplus$ is defined as the multiplication of probabilities. The main assumption underlying the DAMP operations is that the input Distributions are mutually exclusive and independent. This assumption allows DAMP to compute probabilities extremely efficiently, as the operations are simple and can be easily vectorized.
Differentiable Top-$k$ Proofs. Differentiable Top-$k$ Proofs (DTKP) (Huang et al., 2021) was proposed to overcome the shortcomings of DAMP. This provenance tracks a set of up to $k$ proofs for each symbol. Each proof, in turn, denotes the set of input symbols necessary to derive the output symbol. These proofs are then used to compute the probabilities of the output symbols.
In Scallop, DTKP tags are converted into probabilities via differentiable weighted model counting (WMC). This form of DTKP is computationally hard and is by nature difficult to vectorize due to the varying sizes of proof sets and the WMC procedure. We therefore design a vectorized version of DTKP, called DTKP-AM (DTKP with Add-Mult), that can be efficiently computed on GPUs.
We first define the structure of tags in DTKP-AM in a manner that conforms to the constraints of PyTorch tensors. Each tag $t$ for a symbol $s$ is a 2-dimensional tensor of shape $(k,|I|)$, where $k$ is the maximum number of proofs to be retained and $I$ is an ordered list of all input symbols (symbols that are present in the input Distributions). Each row $t_{i}$ of $t$ corresponds to one of the tag’s $k$ proofs. Each element $t_{ij}$ thus represents the probability of the $j$th input symbol in the $i$th proof:
$t_{ij}=\begin{cases}p_{j}&\text{if the $j$th symbol is present in the $i$th proof}\\ \mathbf{\hat{0}}_{ij}&\text{otherwise}\end{cases}$
where $p_{j}$ is the probability of the $j$th input symbol. The probability of each proof is then computed by taking the product of the normalized elements of that proof:
$\Pr(t_{i})=\prod_{j}\text{norm}(t_{ij})\quad\text{where}\quad\text{norm}(t_{ij})=\begin{cases}1&t_{ij}=+\infty\\ 0&t_{ij}=-\infty\\ t_{ij}&\text{otherwise}\end{cases}$
We next define the operations of DTKP-AM in Table 1. The $\oplus$ operation is defined as the union of two tag tensors $t$ and $t^{\prime}$ while $\otimes$ is defined as the element-wise minimum of the absolute values of all possible combinations of proofs in $t$ and $t^{\prime}$. In each case, the $\text{top}_{k}$ operation retains only upto $k$ proofs with the highest probabilities.
These definitions thus allow us to take advantage of the benefits of the DTKP provenance while enabling efficient computation on GPUs. To calculate the probability of the entire tag, DTKP-AM adds the probabilities of the individual proofs and clamps it at 1.
3.4 Building the Dolphin Program
The programmer specifies the neurosymbolic task using a Python program $P$ that uses Dolphin’s programming interface to connect the neural components (e.g., neural networks) with the symbolic components and operations. We call $P$ the symbolic program. Because $P$ is a Python program and Dolphin interfaces with PyTorch, Dolphin supports any PyTorch-based neural network(s), most Python language features, and custom user-defined functions. This feature enables greater flexibility and expressiveness in neurosymbolic programs than existing frameworks.
In addition to $P$, the programmer provides one or more neural networks $M_{1},\ldots,M_{k}$, and a dataset $\mathcal{D}$. Given these inputs, Dolphin extracts the computation graph that encodes how the neural network outputs are transformed using symbolic operations to produce a final result $D_{\text{res}}$. All computations in Dolphin are expressed using distribution objects $D_{i}$. Each Dolphin primitive (e.g., Apply) takes one or more distribution objects as inputs and applies a transformation to produce another distribution object. Because each distribution object $D_{i}$ only contains vectors of tags (or probabilities), the entire computation graph (including the neural network(s)) can be ported to a GPU for faster execution. During training, Dolphin optimizes over the standard objective function:
$\phi(\theta)=\min_{\theta}\sum_{(x,y)\in\mathcal{D}}\mathcal{L}(\text{$P$}(M_{\theta}(x)),y)$ (4)
Here $\mathcal{L}$ is the loss function, such as binary cross entropy. While Dolphin allows $P$ to take multiple neural networks as inputs, we show only one neural network model $M$ here for simplicity.
4 Experiments
We evaluate Dolphin on a set of 13 benchmarks of varying complexity and scale across 5 neurosymbolic tasks. Our evaluation addresses the following research questions:
  • RQ1: Scalability. How does Dolphin scale to complex problems and large datasets compared to existing baselines?
  • RQ2: Accuracy. Do models written in Dolphin converge to SOTA accuracies?
  • RQ3: Provenance Comparisons. Which provenances are most effective for each benchmark?
4.1 Benchmarks
MNIST-SumN. The MNIST-SumN (or briefly, SumN) task from (De Smet et al., 2024) takes as inputs $N$ handwritten digits from the MNIST dataset and returns their sum. We consider three versions of this task: small ($N=5$), medium ($N=10$), and large ($N=15$). The perception model for this task is a standard CNN while the symbolic program sums the outputs of this model.
Hand-Written Formula (HWF). The HWF task from Li et al. (2020) takes as input a set of images of handwritten digits and arithmetic operators representing a formula. The task is to evaluate the formula and return the result. We consider three versions of HWF: small (formulas of length up to 7), medium (formulas of length up to 15), and large (formulas of length up to 19). Similar to MNIST-SumN, the perception model for this task is a standard CNN. The Dolphin program for this task builds strings of formulae from the outputs of the neural model and evaluates them using Python’s eval function, demonstrating the ability of Dolphin to support black-box functions.
PathFinder. PathFinder (or Path) from  Tay et al. (2020) tests the ability of an agent to reason over long-range dependencies within an input image. The image consisting of two dots and a sequence of curved and dashed lines. The task is to identify whether the two dots are connected via the lines. We consider three versions of this task based on the image size in pixels: small (32 x 32), medium (128 x 128), and large (256 x 256). The perception model for this task is also a standard CNN. However, the Dolphin program for this task is recursive, since it must search for paths between the two dots.
CLUTRR. In this task from Sinha et al. (2019), the input is a passage of text containing some information about several individuals and some of their relationships. The task is then to infer the relationship between two given individuals, which is not explicitly provided in the input. Hence, the Dolphin program for this task needs to reason over long chains of relationships recursively. We consider two versions of this task, where the training data contains relation chains of lengths up to 3 (small) or 4 (medium). We use Roberta-base (Liu, 2019) as the language model for this task.
Mugen. In this task from Hayes et al. (2022), the input is a video of gameplay footage that is 3.2 seconds long and a natural language passage captioning the video. The goal is to measure how aligned the text is with the video. This task therefore has two variants: Mugen-TVR, where the model must retrieve the video that best aligns with the text, and Mugen-VTR, where the model must retrieve the text that best aligns with the video. This benchmark tests the ability of the model to reason over multimodal data. We consider two versions of this task: small, with 1000 training samples, and medium, with 5000 training samples. We use a combination of DistilBert (Sanh, 2019) and S3D (De Smet et al., 2024) as the perception model for the text and video inputs respectively.
4.2 Experimental Setup and Baselines
Setup. All experiments, with the exception of the CLUTRR benchmark, were run on machines with two 20-core Intel Xeon Gold 6248 CPUs, four NVIDIA GeForce RTX 4352 GPUs, and 768 GB RAM. Since the CLUTRR benchmark requires a larger amount of GPU memory, it was run on a machine with 8 NVIDIA A100 40GB GPUs instead. We ran each tool thrice until convergence or until a timeout of 10 hours was reached and report the average best accuracy and training time. The code of Dolphin and the benchmarks is provided in the supplementary material.
Baselines. We select Scallop (Li et al., 2023), a contemporary state-of-the-art neurosymbolic framework supporting differentiable programming optimized to run on the CPU and use multiple cores to parallelize its computations. We also choose two sampling-based gradient approximation methods, ISED (Solko-Breslin et al., 2024) and IndeCateR+ (De Smet et al., 2024). We compare Dolphin against Scallop on all benchmarks, and against ISED and IndeCateR on MNIST-SumN and HWF.
4.3 RQ1: Scalability
Task Dolphin Scallop ISED IndeCateR+
$\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$
SumN (S) 78.33 15.67 923.78 184.76 11.80 299.63 59.93 3.82 3.16e3 632.52 40.36
SumN (M) 144.92 14.49 3.41e3 341.57 23.57 2.16e3 216.54 14.94 2.63e4 2.63e3 181.88
SumN (L) 220.47 14.70 7.41e3 493.87 33.60 9.8e3 653.39 44.45 TO 2.27e3 154.37
HWF (S) 3.17e3 158.79 9.99e3 499.57 3.15 1.58e4 790.04 4.97 1.76e4 877.92 5.52
HWF (M) 1.46e4 731.67 TO 1.49e4 20.41 TO 6.83e3 9.34 TO 3.6e3 4.93
HWF (L) 2.42e4 1.21e3 TO 3.92e5 323.21 TO 1.05e4 8.61 TO 5.59e3 4.61
Path (S) 1.08e4 1.08e3 2.2e4 2.2e3 2.03 N.A.
Path (M) 1.79e4 1.79e3 TO 4.17e3 2.32
Path (L) 1.94e4 1.94e3 TO 1.12e4 5.81
CLUTRR (S) 1.54e3 154.85 4.29e3 429.97 2.77 N.A.
CLUTRR (M) 2.91e3 291.36 7.83e3 783.11 2.69
Mugen (S) 3.62e3 180.80 1.34e4 133.68 0.74 N.A.
Mugen (M) 1.78e3 890.34 TO 634.86 0.71
Table 2: Comparison of training times taken by each baseline. The Timeout (TO) is set at 10 hours. $\alpha$ is the scaling factor, which is the ratio of the per epoch training times of the baselines and Dolphin.
Table 2 presents the training times for Dolphin and baselines on all benchmarks, along with the total training time ($T_{\textit{total}}$), time per epoch ($T_{\textit{epoch}}$), and the scaling factor $\alpha$. The scaling factor is the ratio of the per epoch times of the baselines to Dolphin. We observe that in almost all cases, Dolphin shows significant improvements in training times, training up to about 300x faster than Scallop, 44x faster than ISED, and 180x faster than IndeCateR+. On average, Dolphin reports a speedup of 36x times over all baselines. Further, Dolphin converges on all benchmarks, while the baselines time out on most of the Medium and Large versions of the benchmarks. These results thus demonstrate that, unlike existing tools, Dolphin can scale to complex problems and large datasets.
Dolphin trains slightly slower than Scallop on both versions of the Mugen benchmark. This is because the Dolphin program written for Mugen uses Python objects and operations that are not fully batchable across samples. In contrast, the Scallop program, which is written in a compiled and optimized language, runs around 1.3x faster than Dolphin on average per iteration. However, Dolphin requires only 20 epochs to converge whereas Scallop requires almost 1000 epochs (Li et al., 2023). As a result, Dolphin’s total training time is still significantly lower than Scallop’s (3̃.7x for Mugen(S)). We show the accuracy curves for Mugen in Appendix B.
4.4 RQ2: Accuracy
Figure 4 presents the accuracy of Dolphin and the baselines on the different benchmarks. Dolphin accuracies are marked in blue. In all cases, for Dolphin, we report the accuracies of the best-performing provenance. We use the DAMP provenance for MNIST, CLUTRR, and Mugen benchmarks, and the DTKP-AM provenance for the HWF and PathFinder benchmarks.
We observe that in all cases, Dolphin achieves state-of-the-art accuracy, except in CLUTRR, where Dolphin’s accuracy is slightly lower than that of Scallop. For most Medium and Large versions of the benchmarks, Dolphin achieves better accuracy, whereas the baselines either report lower accuracy due to the complexity of the benchmark (e.g., HWF) or fail to converge within the time limit (e.g., Scallop on PathFinder-Large). Most importantly, these results show that Dolphin’s scalability improvements do not come at the cost of accuracy.
Figure 4: Accuracy of Dolphin and baselines across all benchmarks.
4.5 RQ3: Provenance Comparisons
Figure 5: Accuracy and average training time per epoch for DAMP and DTKP-AM.
We perform ablation studies to compare the effectiveness of the DAMP and DTKP-AM provenances for each benchmark. Figure 5 presents the accuracy of Dolphin using DAMP and DTKP-AM in the top graph, and the average training time per epoch in the bottom. We observe that the relative effectiveness of the provenances differ across benchmarks. For instance, in the MNIST-SumN benchmark, the DAMP provenance is more effective than the DTKP-AM provenance, since the top-k proofs are not able to capture all the possible ways in which the sum of the digits can be computed. In contrast, for HWF, the DTKP-AM provenance is more effective than DAMP since the tags in DAMP provenance lack the structure needed to capture the semantics of the symbolic program. Otherwise, DTKP-AM mostly has a slight edge over DAMP. On the other hand, DAMP is faster than DTKP-AM for all benchmarks, since it does not track individual proofs. We thus see that selecting provenances is a trade-off between accuracy and training time.
5 Related Work
Neurosymbolic programming frameworks. Frameworks like Scallop (Li et al., 2023), DeepProbLog (Manhaeve et al., 2018), and ISED (Solko-Breslin et al., 2024) provide a simple interface for neurosymbolic programming. There are also domain-specific tools like NeurASP Yang et al. (2023) for answer set programming and NeuralLog Chen et al. (2021) for phrase alignment in NLP. While these frameworks provide intuitive abstractions, they are bottle-necked due to expensive data transfers between symbolic computations done on CPU versus neural computations that execute on GPU, making neurosymbolic learning hard to scale. In contrast, Dolphin provides a deeper integration of the two worlds by building a Python-based API on top of PyTorch, which scales better.
Scalability improving techniques. Several optimization techniques have been proposed to improve the scalability of differentiable reasoning algorithms. For instance, Greedy NTP (Minervini et al., 2020a) reduces the computation cost of Neural Theorem Prover (NTP) (Rocktäschel & Riedel, 2017) by tracking only a subset of proof states using nearest neighbor search. Likewise, the conditional theorem prover (Minervini et al., 2020b) employs a machine learning-based proof selection technique. However, unlike Dolphin, these methods are point solutions that do not fundamentally address the scalability challenge for neurosymbolic learning.
Specialized neurosymbolic solutions. There are many specialized solutions for various neurosymbolic tasks. For instance, NGS (Li et al., 2020) uses a hand-coded syntax to specify the structure of mathematical expressions for HWF. More general solutions, such as NS-CL (Mao et al., 2019) includes a framework for visual question answering that learns symbolic representations for text and images. NeRd (Chen et al., 2021) transforms questions in natural language into executable programs based on symbolic information extracted from text. In this work, we rather focus on building a general programming framework that improves the scalability of diverse neurosymbolic programs.
6 Conclusion and Limitations
We proposed Dolphin, a programmable framework for scaling neurosymbolic learning. Dolphin provides abstractions for writing symbolic programs along with pluggable vectorized provenances to compute symbolic gradients. This allows users to write differentiable symbolic programs in Python within PyTorch pipelines that can scale to complex programs and large datasets. We show that Dolphin scales significantly better than existing neurosymbolic frameworks while achieving state-of-the-art performance on a variety of tasks.
A limitation of Dolphin is that it requires the user to express symbolic programs in a batched manner. While this is a common pattern within deep learning, it may be restrictive to users new to batched programming. Another limitation is that while Dolphin works well with most models, the representation of Distributions needed by generative models (e.g., Causal LLMs like Llama) has not yet been investigated. We leave this to future work.
Appendix A Dolphin Experiment Details for Benchmarks
The following are more details on the experiment setup. For each experimental trial, we report the highest evaluation accuracy over all epochs. Unless otherwise noted, each trial was run on a machine with Intel Xeon Gold 6248 (2.50 GHz) CPUs and NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs.
A.1 MNIST Sum-N
Each of the MNIST Sum-N tasks had a batch train size of 64 samples, a learning rate of 0.001, and a top-k value of 1. Each of the tasks were trained on a dataset size of the original MNIST dataset divided by N of Sum-N. Sum5’s dataset consisted of 12000 train samples and 2000 test samples. Sum10’s dataset consisted of 6000 train samples and 1000 test samples. Sum15’s dataset consisted of 4000 train samples and 666 test samples.
A.2 Hand-Written Formula
For Hand-Written Formula we trained each task with a batch train size of 64 samples. The learning rate was 0.0001, sample-k was 7, and top-k value was 3. Length 7’s dataset consisted of 9600 samples for training, 2400 samples for testing. Length 15 consisted of 24000 training samples and 6000 testing samples. Length 19 consisted of 32000 training samples and 8000 testing samples.
A.3 PathFinder
For each of the PathFinder tasks, we used a batch train size of 64 samples, a learning rate of 0.0001, and a top-k value of 3. Each task’s dataset consisted of 539459 images for training and 59940 images for testing. Each task had its own dataset of images with dimensions of the task’s pixel size.
A.4 CLUTRR
For each CLUTRR task, we used a single A100 GPU (40 GB), with a learning rate of 0.00001 and use a batch size of 16. The length of the training dataset for CLUTRR (Small) was 11,093 and that of the test set was 1146. The training set for CLUTRR (Medium) contained 15,083 samples and the test set contained 1048 samples.
A.5 Mugen
For each Mugen task, we use a batch size of 3 and a learning rate of 0.0001. From the full Mugen dataset, we sample a training set of 5000 examples for Mugen (Medium), and from that set, we sample a training set of 1000 for Mugen (Small). Both Small and Medium are evaluated on a fixed holdout set of 1000 samples. We train and evaluate for up to 100 epochs.
Appendix B Accuracy Curves for Mugen
Figure 6: Test accuracies for each epoch of Dolphin and Scallop for Mugen (Small). The accuracies for TVR are on the left and VTR are on the right.
Figure 7: Test accuracies for each epoch of Dolphin and Scallop for Mugen (Medium). The accuracies for TVR are on the left and VTR are on the right.
Figures 6 and 7 show the test accuracies for each epoch of Dolphin and Scallop for Mugen (Small) and Mugen (Medium) respectively. The accuracies for TVR are on the left and VTR are on the right. Note that Dolphin converges within 20 epochs in all cases, while Scallop does not converge within 100 epochs.